for CH4(g) + 3Cl2(g) ——> CHCl3(g) + 3HCl(g) how many grams of CH4 is needed to produce 45.5 g CHCl3


1
Expert's answer
2021-10-27T06:41:44-0400

Solution:

The molar mass of CH4 is 16.04 g/mol

The molar mass of CHCl3 is 119.38 g/mol


Calculate the moles of CHCl3:

(45.5 g CHCl3) × (1 mol CHCl3 / 119.38 g CHCl3) = 0.38 mol CHCl3


Balanced chemical equation:

CH4(g) + 3Cl2(g) → CHCl3(g) + 3HCl(g)


According to stoichiometry:

1 mol of CH4 produces 1 mol of CHCl3

x mol of CH4 produces 0.38 mol of CHCl3:

Thus,

x = (0.38 mol CHCl3) × (1 mol CH4 / 1 mol CHCl3) = 0.38 mol CH4


Calculate the mass of CH4:

(0.38 mol CH4) × (16.04 g CH4 / 1 mol CH4) = 6.0952 g CH4 = 6.10 g CH4


Answer: 6.10 grams of CH4 is needed.